Circular No. 7939

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 O4 On July 22, a report of an apparent visual comet discovery was received

from Sebastian Hönig, Dossenhiem, Germany. The object clearly showed motion to the north. Due to poor initial positions and bright moon, attempts to recover the object visually by A. Hale (Cloudcroft, NM) and Hönig were unsuccessful. On July 27, S. Nakano (Sumoto, Japan) reported that K. Kadota (Ageo, Saitama, Japan), following a request by A. Nakamura (Kuma, Ehime, Japan), successfully imaged a comet while searching for the object reported on July 22. Additional observations were received after the comet was posted on the NEO Confirmation Page. Selected observations appear below.

2002 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 22.001  23^{h}25^{m}  +32^{o}40′  12  Hönig
27.64622 23 22 22.65 +40 36 26.1  Kadota
27.66133 23 22 20.99 +40 38 06.1 12.1  _{”}
27.83285 23 22 03.47 +40 57 03.1 15.0 Griesser
27.83316 23 22 03.34 +40 57 04.7  Kušnirák
27.83366 23 22 03.29 +40 57 08.8  Toth
27.83745 23 22 02.88 +40 57 32.8 15.1 Busch
27.84500 23 22 02.14 +40 58 23.9 15.1 Apitzsch
27.84763 23 22 01.89 +40 58 39.7 12.0 Sanchez
27.86705 23 21 59.76 +41 00 51.9 15.2 Payer

S. Hönig (Dossenheim, Germany). 0.25-m f/6.3 Schmidt-Cassegrain tele-

scope. Coma diameter 2′. K. Kadota (Ageo, Saitama, Japan). 0.18-m f/5.5 reflector + CCD. Comet

diffuse; coma diameter 1′.5; tail 1′.8 long in p.a. 213^{o}. M. Griesser (Wiesendangen, Switzerland). 0.40-m f/5.9 hypergraph +

CCD. P. Kušnirák (Ondřejov, Czech Republic). 0.65-m f/3.6 reflector + CCD.

Coma diameter 2′; tail 4′ long towards the southwest. J. Toth (Modra, Slovenia). 0.6-m f/5.5 reflector + CCD. M. Busch and F. Hormuth (Heppenhiem, Germany). 0.45-m f/4.4 reflector

+ CCD. Coma diameter 30′′. R. Apitzsch (Wildberg, Germany). 0.24-m f/10 reflector + CCD. Coma

diameter 1′; tail 2′ long towards the southwest. S. Sanchez (Mallorca). 0.40-m f/2 Schmidt + CCD. Coma diameter 1′; tail

3′ long in p.a. 215^{o}. T. Payer (Düsseldorf, Germany). 0.32-m f/5.7 reflector + CCD.

2002 July 27 \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT Carl W. Hergenrother